a triangle has an area of 52 yard squared and a height of 13 yard. find its base
Answer by sachi(548)

a triangle has an area of 52 yard squared and a height of 13 yard. find its base
ans:-
area of a tringle= 1/2*base*height=1/2*base*13=52
so base=52*2/13=8 yard ans